OpenStudy (nwolf):

m=-4(8,9) what does y=

OpenStudy (anonymous):

i think u hv not typed the question properly/correctly....pls check

OpenStudy (nwolf):

it says fins an equation of the line given the slope and containing the given point then it says the equation of the line is y=??

OpenStudy (anonymous):

formula to use here is y-y1 = m(x - x1) here we are given m = -4 and x1 = 8 and y1 = 9 so substituting in formula we get y-9 = -4(x - 8) y - 9 = -4x + 32 y = -4x + 32 + 9 y = -4x + 41

OpenStudy (anonymous):

m={-32y,-36y}